• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Saha

    Log entry time 08:02:05 on September25,2005

    Entry number 155879

    keyword=Shift Summary

    Shift Summary
    
    Midnight:  Run 5344 continuing  IHWP OUT
    5344:  100 k 
    5345:  100 k
    5346:  Q2 Scan HV = 1500 volts
    5347:  Q2 Scan HV = 1600 volts
    5348:  7 k. 
    Controls network crashed. Rebooted the controls network. 
    Target IOC crashed as well.
    The right spectrometer magnet controls did not reset, so did a hard reboot.
    5349: 100 k
    5350: 100 k
    5351: 100 k
    5352: 100 k
    
    IHWP changed to IN
    
    5253: 100 k
    5254:  continuuing to next shift
    
    
    Compton Run: 12292 continuing
    12292: 96 k 
    12293: